# Probing AGN Feedback in Dwarf Galaxies with Spatially Resolved NIR Coronal Lines from JWST

### v1

@@ 3.1 Emission line fitting and spatial maps generation
To fit the various coronal emission lines (Table 2) and generate spatial flux and velocity maps, we used the Bayesian AGN Decomposition Analysis for Sloan Digital Sky Survey (SDSS) Spectra (BADASS v10.2.0, 2021MNRAS.500.2871S) 111https://github.com/remingtonsexton/BADASS3 code. Through Markov Chain Monte Carlo routines, BADASS performs simultaneous multicomponent fits to emission-line spectra. Individual spaxels in the NIRSpec cubes were fit iteratively across the full spatial extent of the data cubes, where a third-order Legendre polynomial was used for the continuum. The host galaxy and the stellar line-of-sight velocity distribution were fit using the penalized pixel fitting method (pPXF, 2004PASP..116..138C) with templates from the eMILES (2016MNRAS.463.3409V) stellar library. Emission lines were modeled using Gaussian profiles. A secondary Gaussian component to the fit was included based on the line-testing option that is implemented within BADASS. We used a combination of multiple line testing options such as CHI2_RATIO and a modified F-test. The CHI2_RATIO test is the ratio of the reduced chi-squared values for each fitted model, and can be interpreted as the fraction of the difference in the residuals weighted by the noise. The modified F-test generates a confidence between 0 and 1 of the Monte-Carlo resampled maximum likelihood values of the simple and complex models. If the confidence is 0.9, the difference between residuals is significant, and an additional component is justified with 90 confidence. If the value of the CHI2_RATIO was 0.1 and if the modified F-test confidence parameter was 0.9, then adding a second component provided a significant improvement to the fit and was thus justifiable. This secondary component was restricted to have a larger velocity offset from the rest-frame wavelength (see Table 2) and a larger dispersion than the primary component (See Figure 2 for an example fit to a single spaxel and also Section 3.3). Due to the diverse emission profiles in our data, we left the amplitude and width as unrestricted free parameters for the primary component. We also set a signal-to-noise ratio (S/N) threshold of three for all lines and components within each spaxel to determine a detection. Uncertainties were derived from the error spectrum and the random errors associated with the Bayesian fitting process.
3.2 Flux, voff and W80 maps
@@
Maps of the narrow component for a sample of CLs detected in our sample can be seen in Figure 3. The voff shows the offset velocity (calculated from the best-fit Gaussian) from the expected central wavelengths from the systemic redshift (calculated from a combination of stellar absorption features and the strong Paschen 1.873 m line) of the associated CL. The central wavelengths of each line were determined from the wavelengths compiled by CLOUDY222https://linelist.pa.uky.edu/atomic/. W80 is the line width defined to encompass 80% of the total flux, such that W80 = v90 - v10, where v10 and v90 are the velocities at the 10th and 90th percentiles of the total flux, respectively, calculated starting from zero intensity on the red side of the line. For a Gaussian, W80 is 1.0833 times the Full Width at Half Maximum (FWHM). Thus, W80 can be used as an indicator of turbulent or disturbed gas in the galaxy or as the distribution of gas velocities in an outflow (2013MNRAS.436.2576L; 2017A&A...606A..36C; 2020A&ARv..28....2V; 2021MNRAS.505.5469S). All reported W80 values have been corrected for instrumental broadening. The flux maps indicate that most of the emission arises from a compact source at the center of each target. While several previous studies have suggested preferentially blue-shifted emission line profiles for CLs (2009MNRAS.397..172G; 2021ApJ...922..155M), we do not observe such a systematic preference. On the contrary, there seems to be a variety of observed voff within the CL, ranging from blue-shifted to red-shifted values.
3.3 Presence of a broad component
@@
Two of the targets (J0954 and J1009) have a significant number (4) of CLs that show signs of having broader wings than a single Gaussian (Figures 2 and 4), as determined by the line-component testing for the spectra in the individual spaxels (Section 3.1). Previous works have indicated that the optical coronal emission line profiles can be fitted by two-Gaussian components (2005MNRAS.364L..28P; 2006ApJ...653.1098R; 2011ApJ...739...69M). These studies also find that the first component extends to large distances (0.7 - 3.8 kpc). The second broader, weaker component cannot be measured outside of the bright nucleus (0.03 kpc). We find a similar trend in the two dwarf galaxies where broad CLs are detected (J0954 and J1009). We also find that the broad component is detected only in the stronger lines of [Si VI], [Si VII], [Mg IV], [Mg VIII], and [Ar VI]. The broad component is too weak to detect in weaker lines, including those with larger I.P. as well as in Ca and Al lines which have similar I.P. as the aforementioned lines. The second component in all targets is significantly broad (W80 300 km s-1), but does not have significant velocity offset. For one of the targets (J0906), a second component was not included based on the line tests, but we find that the narrow component is significantly broad (W80 300 km s-1).
@@
The presence of continuum “wiggles” (Section 2.2) in the data significantly limits the detailed characterization of the broad component. This posed challenges in fitting and getting resolved maps of the broad component. To fit the individual spaxels, we had to model the wiggles and remove them from the data. We found that it was impossible to distinguish between the wiggles and the faint wings of the outflows. To quantify the difference between the wiggled and de-wiggled cubes, we extract and fit spatially integrated spectra from each cube using the same aperture. We find that the parameter that was most affected by the uncertainty introduced by the wiggles is the width of the broad component, whereas fluxes and velocity offsets do not change significantly. We find that the width parameters, such as W80 and FWHM of the broad components, were lower by 20-30 in the de-wiggled cube as compared to the wiggled cubes. Thus all the estimates of the W80 of the broad component in this work should be treated as lower limits.
3.4 Determining the extent of the CLs
@@
The de-projected extent in spaxel units is calculated using Equation (1) from 2021ApJ...920...62N:
We then convert the extent to physical units (kpc) using the astropy.cosmology Python package. We also estimate the extent using alternative definitions, including the half-light radius, 95th-percentile radius (R95) and root-mean-square (rms) radius, and find that the trends described in Section 5.2 remain valid across these methods.
@@
4 Results
4.1 Coronal line detections
@@
Note. — Columns: (1) Coronal line species; (2) Rest wavelength from CLOUDY (m). # indicates wavelength from the CAFE (2025ascl.soft01001D) linelist111https://github.com/GOALS-survey/CAFE. (3) Ionization potential (eV) from the Chianti database222https://www.chiantidatabase.org. (4) Critical density (log cm-3) from CLOUDY. (5–8) Detection status in each galaxy: = detected with JWST/NIRSpec; = broad component detected; † = previously detected with Keck/NIRES; ∗ = detected only in the 0.3 integrated spectrum (not across individual spaxels).
@@
Previous ground based NIR observations using Keck/NIRES (2021ApJ...911...70B) revealed the presence of high-ionization CLs in the four targets between 0.9 - 2.3 m. We now report the detection of new CLs due to the improved sensitivity as well as in the extended wavelength range provided by JWST/NIRSpec (0.9 - 5 m). 2021ApJ...911...70B reported a single CL, [Ca VIII], in J0842, which we do not detect in our JWST/NIRSpec observations. However, the expected wavelength of [Ca VIII] lies very close to the detector gap in the data cube, which makes it challenging to detect robustly. We also detect [Mg IV] at 4.48 m, which is the only coronal line reliably observed in this target with NIRSpec. For J0906, 2021ApJ...911...70B identified five CLs from ground-based data. We recover four of these in the NIRSpec spectra. [Ca VIII] is not detected as it falls within the transmission gap between the two NIRSpec detectors. In addition to the previously known lines, we detect six new CLs, increasing the total number of NIR-detected CLs in this galaxy to 11. In J0954, three NIR CLs had been previously reported (2021ApJ...911...70B). Our observations reveal 11 additional CLs, bringing the total to 14. Finally, J1009 exhibits the richest coronal line spectrum among the four targets. While previous ground-based observations identified four CLs, our NIRSpec data reveal 12 additional lines. These include several highly ionized species such as [S XI] (I.P. = 447 eV), as well as multiple ionization states of Ca and Na. For a complete list of detected CLs in the four targets see Tab. 2 and Figure 5.
@@
Overall, we detect all CLs reported by 2021ApJ...911...70B except those that fall in the JWST detector gap. We do not confirm the [Ca VIII] detection reported by 2021ApJ...911...70B; given its 2.2 significance, it may have been a noise artifact. Among the new lines, only one ([S XI]) lies within the Keck/NIRES wavelength range but was likely missed due to lower sensitivity or imperfect atmospheric subtraction. [Si X] was not detected with Keck/NIRES in J0954 but is seen with NIRSpec, likely due to the higher spectral resolution. The remaining new detections are enabled by the extended wavelength coverage of NIRSpec.
@@
Three of the four targets also show optical CLs in prior SDSS, Keck/LRIS, Keck/KCWI and Gemini/GMOS spectra (2019ApJ...884...54M; 2021ApJ...911...70B). J0906 has [Ne V] 3426, [Fe X] 6374 and [Fe VII] 5721, 5159, 6087. J0954 has [Ne V] 3346, 3426, [Fe V] 3839, and [Fe VII] 5159, 6087. J1009 has possible detections of [Fe VII] 5721, 6087. J0842 has no previously detected optical CLs.
@@
The CL peaks in all targets coincide with the K-band continuum peak, and, within each target, the CL peaks coincide with one another, strongly suggesting that they are produced by the central AGN source. Although CLs have also been attributed to tidal disruption events (TDEs; 2023MNRAS.519.2035H; 2024MNRAS.528.7076C), a comparison between the JWST spectra and Keck/NIRES observations taken 5 years earlier (2021ApJ...911...70B) shows no evidence of fading of the CLs. A five-year interval is consistent with the expected fading timescales of TDE-driven CLs (2013ApJ...774...46Y). Thus, it is unlikely that these CLs are produced by a TDE and are instead directly ionized by the central AGN.
4.2 Source of Ionization
@@
While high-ionization CLs are generally attributed to photoionization by the hard radiation field of the AGN, some studies have proposed that they may also be produced by fast radiative shocks associated with outflows (1995ApJ...455..468D). Given that the dwarf galaxies in our sample exhibit evidence of outflows, we investigate the dominant ionization mechanism responsible for the observed CLs.
@@
To disentangle the contributions from AGN photoionization and shocks, we utilize diagnostic flux ratios such as [Si VI]/Br and [Si VII]/Br vs. [Fe II] 1.64 m/Br. The [Fe II] line is a well-established tracer of shock excitation (2013ApJ...775..115U), whereas Br traces star formation and the overall ionizing photon field. These ratios are compared against both AGN photoionization models from 2004ApJS..153...75G; 2004ApJS..153....9G and shock+precursor models from 2008ApJS..178...20A, as implemented in the IDL Tool for Emission-line Ratio Analysis (ITERA; 2010NewA...15..614G). We plot the ratios obtained from the total fluxes of the emission lines from individual spaxels in two of the galaxies (J0954 and J1009), where all the lines are present in Figure 6.
@@
As shown in Figure 6, most of the spatially resolved emission-line ratios for J0954 and J1009 align more closely with the AGN photoionization models. Within the AGN photoionization grids, the points scatter around the gridline with more negative , indicating a steeper power-law spectrum, which can be linked to factors like higher accretion rates and potentially lower black hole masses (2005A&A...432...15P; 2010A&A...512A..58I). The results are still valid when we use other CLs (i.e. [Si VII]). We assume solar metallicity for the grids, similar to 2020ApJ...905..166L and 2021ApJ...911...70B, but we find that if we assume lower metallicities for these dwarf galaxies (2020ApJ...903...58C), the results are still broadly valid. There is an upward shift of the spaxels towards the shock+precursor grids as well as towards higher and log U values in the AGN grids for lower metallicities, but this shift is 0.1 dex for 0.5Z⊙ and 0.2 dex for 0.25Z⊙.
@@
We further examine how the line ratios change with radial distance from the nucleus for the two targets. In J1009, we do not see any clear trend between the radius and the line ratios. However, for J0954, we find that near the galaxy center the ratios are consistent with pure AGN photoionization. At larger radii, they shift slightly towards values predicted by the shock+precursor models (Figure 7). However, this shift is not strong enough to require a significant contribution from shocks, as the ratios also remain broadly consistent with the photoionization models. Furthermore, since the CL emission follows the AGN ionization bicone (see Section 5.3), declines with radius, and does not show excess emission at larger distances, any contribution from shocks is likely small.
5 Discussion
5.1 Non-Uniform Frequency of the Coronal Lines

### v2

@@ III.1 Emission line fitting and spatial maps generation
To fit the various coronal emission lines (Table 2) and generate spatial flux and velocity maps, we used the Bayesian AGN Decomposition Analysis for Sloan Digital Sky Survey (SDSS) Spectra (BADASS v10.2.0, Sexton et al., 2021) 111https://github.com/remingtonsexton/BADASS3 code. Through Markov Chain Monte Carlo routines, BADASS performs simultaneous multicomponent fits to emission-line spectra. Individual spaxels in the NIRSpec cubes were fit iteratively across the full spatial extent of the data cubes, where a third-order Legendre polynomial was used for the continuum. The host galaxy and the stellar line-of-sight velocity distribution were fit using the penalized pixel fitting method (pPXF, Cappellari and Emsellem 2004) with templates from the eMILES (Vazdekis et al., 2016) stellar library. Emission lines were modeled using Gaussian profiles. A secondary Gaussian component to the fit was included based on the line-testing option that is implemented within BADASS. We used a combination of multiple line testing options such as CHI2_RATIO and a modified F-test. The CHI2_RATIO test is the ratio of the reduced chi-squared values for each fitted model, and can be interpreted as the fraction of the difference in the residuals weighted by the noise. The modified F-test generates a confidence between 0 and 1 of the Monte-Carlo resampled maximum likelihood values of the simple and complex models. If the confidence is 0.9, the difference between residuals is significant, and an additional component is justified with 90 confidence. If the value of the CHI2_RATIO was 0.1 and if the modified F-test confidence parameter was 0.9, then adding a second component provided a significant improvement to the fit and was thus justifiable. This secondary component was restricted to have a larger velocity offset from the rest-frame wavelength (see Table 2) and a larger dispersion than the primary component (See Figure 2 for an example fit to a single spaxel and also Section III.3). Due to the diverse emission profiles in our data, we left the amplitude and width as unrestricted free parameters for the primary component. We also set a signal-to-noise ratio (S/N) threshold of three for all lines and components within each spaxel to determine a detection. Uncertainties were derived from the error spectrum and the random errors associated with the Bayesian fitting process.
III.2 Flux, voff and W80 maps
@@
Maps of the narrow component for a sample of CLs detected in our sample can be seen in Figure 3. The voff shows the offset velocity (calculated from the best-fit Gaussian) from the expected central wavelengths from the systemic redshift (calculated from a combination of stellar absorption features and the strong Paschen 1.873 m line) of the associated CL. The central wavelengths of each line were determined from the wavelengths compiled by CLOUDY222https://linelist.pa.uky.edu/atomic/. W80 is the line width defined to encompass 80% of the total flux, such that W80 = v90 - v10, where v10 and v90 are the velocities at the 10th and 90th percentiles of the total flux, respectively, calculated starting from zero intensity on the red side of the line. For a Gaussian, W80 is 1.0833 times the Full Width at Half Maximum (FWHM). Thus, W80 can be used as an indicator of turbulent or disturbed gas in the galaxy or as the distribution of gas velocities in an outflow (Liu et al., 2013; Concas et al., 2017; Veilleux et al., 2020; Scholtz et al., 2021). All reported W80 values have been corrected for instrumental broadening. The flux maps indicate that most of the emission arises from a compact source at the center of each target. While several previous studies have suggested preferentially blue-shifted emission line profiles for CLs (Gelbord et al., 2009; Molina et al., 2021), we do not observe such a systematic preference. On the contrary, there seems to be a variety of observed voff within the CL, ranging from blue-shifted to red-shifted values.
III.3 Presence of a broad component
@@
Two of the targets (J0954 and J1009) have a significant number (4) of CLs that show signs of having broader wings than a single Gaussian (Figures 2 and 4), as determined by the line-component testing for the spectra in the individual spaxels (Section III.1). Previous works have indicated that the optical coronal emission line profiles can be fitted by two-Gaussian components (Prieto et al., 2005; Rodríguez-Ardila et al., 2006; Müller-Sánchez et al., 2011). These studies also find that the first component extends to large distances (0.7 - 3.8 kpc). The second broader, weaker component cannot be measured outside of the bright nucleus (0.03 kpc). We find a similar trend in the two dwarf galaxies where broad CLs are detected (J0954 and J1009). We also find that the broad component is detected only in the stronger lines of [Si VI], [Si VII], [Mg IV], [Mg VIII], and [Ar VI]. The broad component is too weak to detect in weaker lines, including those with larger I.P. as well as in Ca and Al lines which have similar I.P. as the aforementioned lines. The second component in all targets is significantly broad (W80 300 km s-1), but does not have significant velocity offset. For one of the targets (J0906), a second component was not included based on the line tests, but we find that the narrow component is significantly broad (W80 300 km s-1).
@@
The presence of continuum “wiggles” (Section II.2) in the data significantly limits the detailed characterization of the broad component. This posed challenges in fitting and getting resolved maps of the broad component. To fit the individual spaxels, we had to model the wiggles and remove them from the data. We found that it was impossible to distinguish between the wiggles and the faint wings of the outflows. To quantify the difference between the wiggled and de-wiggled cubes, we extract and fit spatially integrated spectra from each cube using the same aperture. We find that the parameter that was most affected by the uncertainty introduced by the wiggles is the width of the broad component, whereas fluxes and velocity offsets do not change significantly. We find that the width parameters, such as W80 and FWHM of the broad components, were lower by 20-30 in the de-wiggled cube as compared to the wiggled cubes. Thus all the estimates of the W80 of the broad component in this work should be treated as lower limits.
III.4 Determining the extent of the CLs
@@
The de-projected extent in spaxel units is calculated using Equation (1) from Negus et al. 2021:
We then convert the extent to physical units (kpc) using the astropy.cosmology Python package. We also estimate the extent using alternative definitions, including the half-light radius, 95th-percentile radius (R95) and root-mean-square (rms) radius, and find that the trends described in Section V.2 remain valid across these methods.
@@
IV Results
IV.1 Coronal line detections
@@
Note. — Columns: (1) Coronal line species; (2) Rest wavelength from CLOUDY (m). # indicates wavelength from the CAFE (Diaz-Santos et al., 2025) linelist111https://github.com/GOALS-survey/CAFE. (3) Ionization potential (eV) from the Chianti database222https://www.chiantidatabase.org. (4) Critical density (log cm-3) from CLOUDY. (5–8) Detection status in each galaxy: = detected with JWST/NIRSpec; = broad component detected; † = previously detected with Keck/NIRES; ∗ = detected only in the 0.3 integrated spectrum (not across individual spaxels).
@@
Previous ground based NIR observations using Keck/NIRES (Bohn et al., 2021) revealed the presence of high-ionization CLs in the four targets between 0.9 - 2.3 m. We now report the detection of new CLs due to the improved sensitivity as well as in the extended wavelength range provided by JWST/NIRSpec (0.9 - 5 m). Bohn et al. 2021 reported a single CL, [Ca VIII], in J0842, which we do not detect in our JWST/NIRSpec observations. However, the expected wavelength of [Ca VIII] lies very close to the detector gap in the data cube, which makes it challenging to detect robustly. We also detect [Mg IV] at 4.48 m, which is the only coronal line reliably observed in this target with NIRSpec. For J0906, Bohn et al. 2021 identified five CLs from ground-based data. We recover four of these in the NIRSpec spectra. [Ca VIII] is not detected as it falls within the transmission gap between the two NIRSpec detectors. In addition to the previously known lines, we detect six new CLs, increasing the total number of NIR-detected CLs in this galaxy to 11. In J0954, three NIR CLs had been previously reported (Bohn et al., 2021). Our observations reveal 11 additional CLs, bringing the total to 14. Finally, J1009 exhibits the richest coronal line spectrum among the four targets. While previous ground-based observations identified four CLs, our NIRSpec data reveal 12 additional lines. These include several highly ionized species such as [S XI] (I.P. = 447 eV), as well as multiple ionization states of Ca and Na. For a complete list of detected CLs in the four targets see Tab. 2 and Figure 5.
@@
Overall, we detect all CLs reported by Bohn et al. 2021 except those that fall in the JWST detector gap. We do not confirm the [Ca VIII] detection reported by Bohn et al. 2021; given its 2.2 significance, it may have been a noise artifact. Among the new lines, only one ([S XI]) lies within the Keck/NIRES wavelength range but was likely missed due to lower sensitivity or imperfect atmospheric subtraction. [Si X] was not detected with Keck/NIRES in J0954 but is seen with NIRSpec, likely due to the higher spectral resolution. The remaining new detections are enabled by the extended wavelength coverage of NIRSpec.
@@
Three of the four targets also show optical CLs in prior SDSS, Keck/LRIS, Keck/KCWI and Gemini/GMOS spectra (Manzano-King et al., 2019; Bohn et al., 2021). J0906 has [Ne V] 3426, [Fe X] 6374 and [Fe VII] 5721, 5159, 6087. J0954 has [Ne V] 3346, 3426, [Fe V] 3839, and [Fe VII] 5159, 6087. J1009 has possible detections of [Fe VII] 5721, 6087. J0842 has no previously detected optical CLs.
@@
The CL peaks in all targets coincide with the K-band continuum peak, and, within each target, the CL peaks coincide with one another, strongly suggesting that they are produced by the central AGN source. Although CLs have also been attributed to tidal disruption events (TDEs; Hinkle et al. 2023; Clark et al. 2024), a comparison between the JWST spectra and Keck/NIRES observations taken 5 years earlier (Bohn et al., 2021) shows no evidence of fading of the CLs. A five-year interval is consistent with the expected fading timescales of TDE-driven CLs (Yang et al., 2013). Thus, it is unlikely that these CLs are produced by a TDE and are instead directly ionized by the central AGN.
IV.2 Source of Ionization
@@
While high-ionization CLs are generally attributed to photoionization by the hard radiation field of the AGN, some studies have proposed that they may also be produced by fast radiative shocks associated with outflows (Dopita and Sutherland, 1995). Given that the dwarf galaxies in our sample exhibit evidence of outflows, we investigate the dominant ionization mechanism responsible for the observed CLs.
@@
To disentangle the contributions from AGN photoionization and shocks, we utilize diagnostic flux ratios such as [Si VI]/Br and [Si VII]/Br vs. [Fe II] 1.64 m/Br. The [Fe II] line is a well-established tracer of shock excitation (U et al., 2013), whereas Br traces star formation and the overall ionizing photon field. These ratios are compared against both AGN photoionization models from Groves et al. 2004b, a and shock+precursor models from Allen et al. 2008, as implemented in the IDL Tool for Emission-line Ratio Analysis (ITERA; Groves and Allen 2010). We plot the ratios obtained from the total fluxes of the emission lines from individual spaxels in two of the galaxies (J0954 and J1009), where all the lines are present in Figure 6.
@@
As shown in Figure 6, most of the spatially resolved emission-line ratios for J0954 and J1009 align more closely with the AGN photoionization models. Within the AGN photoionization grids, the points scatter around the gridline with more negative , indicating a steeper power-law spectrum, which can be linked to factors like higher accretion rates and potentially lower black hole masses (Piconcelli et al., 2005; Ishibashi and Courvoisier, 2010). The results are still valid when we use other CLs (i.e. [Si VII]). We assume solar metallicity for the grids, similar to Liu et al. 2020 and Bohn et al. 2021, but we find that if we assume lower metallicities for these dwarf galaxies (Cai et al., 2020), the results are still broadly valid. There is an upward shift of the spaxels towards the shock+precursor grids as well as towards higher and log U values in the AGN grids for lower metallicities, but this shift is 0.1 dex for 0.5 Z⊙ and 0.2 dex for 0.25 Z⊙.
@@
In J0954, a larger fraction of the flux ratios overlap with the shock+precursor model predictions. This fraction likely goes up if we consider lower metallicities. To further examine this, we plot the line ratios as a function of radial distance from the nucleus. We find that near the galaxy center the ratios are consistent with pure AGN photoionization, even while assuming lower metallicities. At larger radii, they shift towards values predicted by the shock+precursor models (Figure 7). This spatial trend suggests that, while AGN photoionization dominates near the nucleus, shocks driven by outflows may contribute to line excitation at larger distances from the central source, potentially accounting for the extended nature of some CLs. These findings are consistent with previous studies in massive galaxies that proposed a composite ionization scenario involving both AGN and shocks (Rodríguez-Ardila et al., 2006). Furthermore, since the CL emission follows the AGN ionization bicone (see Section V.3), declines with radius, and does not show excess emission at larger distances, any contribution from shocks is likely small.
V Discussion
V.1 Non-Uniform Frequency of the Coronal Lines